OpenStudy (anonymous):

find exact value of tan(sin^-1 1/5)

OpenStudy (anonymous):

tan(sin^-1 1/5) = tan ( 11.537 ) = 0.204

OpenStudy (dumbcow):

same as last problem convert tan into expression with sine \[\tan = \frac{\sin}{\cos} = \frac{\sin}{\sqrt{1-\sin^2}}\]

OpenStudy (dumbcow):